OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions.Through (−1, −2); perpendicular to the line 2x + 5y + 5 = 0

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

-5x+2y-1=0

OpenStudy (owlcoffee):

i'll go slow, this kind of problems require a lot of reasoning. we know this: \[m _{1}=-\frac{ 1 }{ m _{2} }\] All this says is, we need a slope, with inverted number in order for them to be perpendicular. Moving on to the problem: we have the ecuation: \[2x+5y+5=0\] Now, we want it in the form y=mx+b, so we can see the slope. \[2x+5=-5y\]= \[\frac{ 2x+5 }{ -5 }=y\] ordering the terms: \[y=-\frac{ 2x }{ 5 }-1\] so, we can see that the slope of the line1 is -2/5 so the slope of the second one is 5/2 we have one of the points it intersects, so we'll use the ecuation: \[y-y _{1}=m(x-x _{1})\] Let's replace the terms: \[y+2=\frac{ 5 }{ 2 }(x+1)\] So the ecuation of the line that satisfies that conditions is: \[y=\frac{ 5 }{ 2 }x+\frac{ 1 }{ 2 }\]